### v1

# The Too Visible QCD Axion Luca Di Luzioa, Michele Redib, Alessandro Strumiac, Andrea Tesib, Arsenii V. Titovd,a

Abstract
Murayama [1] proposed a GeV-scale axion theory where the up-quark mass term is generated dynamically by the QCD chiral condensate, spontaneously breaking a Peccei-Quinn symmetry. It predicts a too large mass splitting between neutral and charged pions. Trying to solve this problem we explore extensions. Despite some partial improvements, we identify a structural obstruction: the new Peccei-Quinn spurion breaks the accidental isospin symmetry of the chiral Lagrangian, leading to an enhanced higher-order operator. As a consequence, pion scatterings too are distorted. We also examine the limit where the axion becomes light, finding that it is excluded by fifth-force constraints.
1 Introduction
In a recent paper Murayama proposed a novel axion theory to address the strong CP problem [1], where all the relevant dynamics takes place at the QCD scale. As well known, QCD solves the strong CP problem if the bare mass of the up quark vanishes, $m_{u}=0$. In this case the $\eta^{\prime}$ acts as an axion, dynamically relaxing the $\theta$-angle to zero. However, this possibility is excluded by comparing lattice computations with data [2]. Building on the massless up quark solution, Murayama added a new complex scalar Standard Model (SM) singlet $\varphi$ coupled to up quarks with unit Peccei-Quinn (PQ) charge under the rotation of $u_{R}$. When quarks develop their condensate, $\varphi$ acquires a PQ-breaking vacuum expectation value effectively giving a mass to the up quark. The main issue is that the $\pi^{0}$ mass is modified
$\frac{\Delta m_{\pi^{0}}}{m_{\pi^{0}}}\approx-\frac{m_{u}}{2(m_{u}+m_{d})}\approx-0.15.$ (1)
The observed pion mass splitting $m_{\pi^{\pm}}-m_{\pi^{0}}\approx 0.03m_{\pi^{0}}$ arises predominantly from electromagnetic self-energy effects, with only a small correction from the up-down quark mass difference. Lattice computations claim a per-cent level precision on the electromagnetic effect [3, 4, 5]. So this model is excluded unless the correction to the neutral pion mass can be nearly totally canceled.
We try improving on this situation. In section 2 we review and extend the model. In section 3 we find some improvement, but only partial. Neither the Kaplan-Manohar ambiguity nor one-loop corrections to pion masses avoid the problem. The pion mass splitting is only mildly reduced by assuming that the down quark $d_{R}$ too is PQ charged. In view of $m_{d}/m_{u}\approx 2$, the pion mass splitting could be avoided by assuming PQ charges $n_{u}/n_{d}=2$; however the new physics gets non-perturbative below the QCD scale. This can be cured by adding new light PQ scalars, but scans in the larger parameter space do not find an allowed solution. Isospin is recovered in a model where the $s$ quark only is PQ charged, but the new PQ scalars turn out to be too light.
In section 4 we identify a structural obstacle. Like quark masses, the interactions of $\varphi$ break the accidental SU(3)u,d,s flavour symmetry of massless QCD. As a result, integrating out QCD and $\varphi$, gives an effective chiral Lagrangian for mesons $U$ that differs from the usual QCD chiral Lagrangian. The operator related to quark masses is substituted by a different higher-order operator that needs to have comparable size,
$\Delta\mathscr{L}_{\rm PQ}\sim\operatorname{tr}[I_{\rm PQ}^{\dagger}U]\operatorname{tr}[I_{\rm PQ}U^{\dagger}]$ (2)
containing the new PQ spurion $I_{\rm PQ}$, given by $I_{\rm PQ}\propto\mathrm{diag}(m_{u}\,,0\,,0)$ in the model of [1]. This spurion breaks the SU(2)u,d isospin symmetry in a higher-order chiral operator, so pion masses get the order one correction of eq. (1), and pion scatterings too are similarly affected. The general problem of the new theories is that they unavoidably predict order unity corrections to meson physics, but the QCD chiral Lagrangian has been more precisely tested.
In the final section 5 we reconsider the new theory abandoning the attempt of a QCD-scale axion. In some limit one recovers the ordinary light axion pseudo-scalar, but it is accompanied by a light scalar, excluded by fifth-force constraints.
2 The model and its generalizations
We start our discussion reviewing and extending the results in [1].
2.1 General PQ charges
We generalise the theory allowing for general PQ charges $n_{q}$ and couplings of $\varphi$ to light quarks $q$, including down and strange quarks. The effective Lagrangian before QCD confinement reads
$\begin{split}\mathscr{L}&=-\frac{1}{4}G^{a}_{\mu\nu}G^{a,\mu\nu}+\theta\frac{g_{s}^{2}}{32\pi^{2}}G^{a}_{\mu\nu}\tilde{G}^{a,\mu\nu}+f_{\varphi}^{2}\bigg[|\partial_{\mu}\varphi|^{2}-m^{2}_{\varphi}|\varphi|^{2}-\lambda f_{\varphi}^{2}|\varphi|^{4}\bigg]+\\ &-\sum_{q=\{u,d,s\}}\bigg(m_{q}\delta_{n_{q}0}+\frac{f_{\varphi}\kappa_{q}}{\sqrt{2}}(\sqrt{2}\varphi)^{n_{q}}\bigg)\bar{q}_{L}q_{R}+\text{h.c.}\end{split}$ (3)
[…]
$\frac{E}{N}\equiv\frac{2}{3}\frac{(4n_{u}+n_{d}+n_{s})}{\sum_{q}n_{q}}\,,\quad\quad N_{\rm DW}=2N=\bigg|\sum_{q}n_{q}\bigg|\,.$ (4)
PQ gets spontaneously broken by the QCD chiral condensate after confinement. On the minimum, the quarks charged under PQ acquire non-vanishing masses
$m_{q}=\kappa_{q}\frac{f_{\varphi}}{\sqrt{2}}\,.$ (5)
The quark couplings can then be rewritten as
$\sum_{q=\{u,d,s\}}m_{q}(\sqrt{2}\varphi)^{n_{q}}\bar{q}_{L}q_{R}+\hbox{h.c.}$ (6)
for both $n_{q}=0$ and $n_{q}\neq 0$.
[…]
$\begin{split}\mathscr{L}_{\pi\varphi}&=f_{\varphi}^{2}\bigg[|\partial_{\mu}\varphi|^{2}-m^{2}_{\varphi}|\varphi|^{2}-\lambda f_{\varphi}^{2}|\varphi|^{4}\bigg]\\ &+\frac{f_{\pi}^{2}}{4}\operatorname{tr}[D_{\mu}UD^{\mu}U^{\dagger}]+\frac{f_{\pi}^{2}B_{0}}{2}\operatorname{tr}[M_{q}U^{\dagger}+M_{q}^{\dagger}U]-\frac{m_{1}^{2}f_{\pi}^{2}}{12}\left[\left(-\frac{i}{2}\ln\det U+\hbox{h.c.}\right)-\theta\right]^{2}.\end{split}$ (7)
This Lagrangian captures the low-energy behaviour of the full theory of eq. (3). The quark mass matrix
$M_{q}=\mathrm{diag}\left[m_{u}(\sqrt{2}\varphi)^{n_{u}},\,m_{d}(\sqrt{2}\varphi)^{n_{d}},\,m_{s}(\sqrt{2}\varphi)^{n_{s}}\right]$ (8)
[…]
$\frac{\Pi}{f_{\pi}}=\begin{pmatrix}\displaystyle\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta}{\sqrt{6}}+\frac{\eta^{\prime}}{\sqrt{3}}&\pi^{+}&K^{+}\\[6.0pt] \pi^{-}&\displaystyle-\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta}{\sqrt{6}}+\frac{\eta^{\prime}}{\sqrt{3}}&K^{0}\\[6.0pt] K^{-}&\bar{K}^{0}&\displaystyle-\frac{2\,\eta}{\sqrt{6}}+\frac{\eta^{\prime}}{\sqrt{3}}\end{pmatrix}$ (9)
in terms of dimension-less fields related to the canonically normalized fields by $\pi^{0}_{\mathrm{can}}=f_{\pi}\pi^{0}$, etc.
The mass spectrum of the model is then obtained by minimizing the potential terms of eq. (7). The novelty is the presence of the new complex scalar $\varphi$, that can be expanded in CP-even, $\sigma$, and CP-odd, $a$, components as
$\varphi=\frac{1+\sigma}{\sqrt{2}}e^{ia}.$ (10)
$\sigma$ is a radial mode, and $a$ is the angular mode related to the axion. The PQ symmetry allows to remove the $\theta$ angle and take the vacuum expectation values of the neutral mesons to be zero. The minimum condition comes from requiring $\sigma=0$ and $a=0$ on the vacuum. The value of $m_{\varphi}$ needed to get $\sigma=0$ is
$m_{\varphi}^{2}=\frac{B_{0}f_{\pi}^{2}}{f_{\varphi}^{2}}\sum_{q}m_{q}n_{q}-\lambda f_{\varphi}^{2}.$ (11)
Since we have assumed that both $m_{\varphi}^{2}$ and $\lambda$ are positive, [^2] If the squared mass term is negative, PQ breaks spontaneously irrespectively of the QCD dynamics, and the model reduces to a DFSZ-like axion model [8, 9]. this model has an upper bound on $f_{\varphi}$. For example, focussing on the model $\{n_{u},n_{d},n_{s}\}=\{1,0,0\}$, we get
$f_{\varphi}\lesssim\text{GeV}\,\left(\frac{5.2~10^{-5}}{\lambda}\right)^{1/4}\,.$ (12)
A large decay constant can only be achieved for a tiny quartic. If $f_{\varphi}$ is well within this upper bound, the minimization condition is independent from $\lambda$, and $m_{\varphi}^{2}$ just depends on the mass of the heaviest quark with PQ charge. In the case of [1], we have [^3] Note that $m_{\varphi}$ is $\sqrt{2}$ times smaller than in [1]. This originates from the normalization of kinetic terms in the chiral Lagrangian.
$m_{\varphi}^{2}=\frac{m_{\pi^{\pm}}^{2}f_{\pi}^{2}}{f_{\varphi}^{2}}\frac{m_{u}}{m_{u}+m_{d}}=(2.3\,\kappa_{u}\,\mathrm{GeV})^{2}.$ (13)
Assuming $\kappa_{u}\sim 1$ implies $f_{\varphi}\approx 3\,\mathrm{MeV}$, while the limit $\kappa_{u},\lambda\sim 0$ realizes a scenario with large $f_{\varphi}$, leading to an invisible axion window discussed in section 5.
The mass of the radial mode is
$m_{\sigma}^{2}=m_{\varphi}^{2}+3\lambda f_{\varphi}^{2}=\frac{B_{0}f_{\pi}^{2}}{f_{\varphi}^{2}}\sum_{q}m_{q}n_{q}+2\lambda f_{\varphi}^{2}\,.$ (14)
We next compute the masses of the CP-odd scalars, $a$ and the mesons $\pi^{0},\eta,\eta^{\prime}$.
2.2 CP-odd scalar mass matrix
Neglecting interactions with the CP-even $\sigma$, which amounts to set $\sigma=0$, the Lagrangian restricted to self-interactions of the CP-odd scalars around the minimum of (11) is
$\begin{split}\mathscr{L}&=\frac{f_{\varphi}^{2}}{2}(\partial a)^{2}+\frac{f_{\pi}^{2}}{2}[(\partial\pi^{0})^{2}+(\partial\eta)^{2}+(\partial\eta^{\prime})^{2}]-\frac{m_{1}^{2}}{2}f_{\pi}^{2}(\eta^{\prime}-\theta)^{2}-B_{0}f_{\pi}^{\,2}\Bigg[m_{s}\cos\!\left(a\,n_{s}+\frac{2\eta-\sqrt{2}\,\eta^{\prime}}{\sqrt{3}}\right)+\\ &+m_{u}\cos\!\left(a\,n_{u}-\frac{\eta}{\sqrt{3}}-\sqrt{\frac{2}{3}}\,\eta^{\prime}-\pi^{0}\right)+m_{d}\cos\!\left(a\,n_{d}-\frac{\eta}{\sqrt{3}}-\sqrt{\frac{2}{3}}\,\eta^{\prime}+\pi^{0}\right)\Bigg].\end{split}$ (15)
It is useful to see explicitly how the strong CP problem is solved in this setup. Under a PQ transformation $\bar{q}_{L}q_{R}\to e^{i\alpha Q_{\rm PQ}}\bar{q}_{L}q_{R}$, with $Q_{\rm PQ}=\text{diag}(n_{u},n_{d},n_{s})$, the pseudo-scalar fields transform as: $a\to a-\alpha$ and $U\to Ue^{-i\alpha Q_{\rm PQ}}$. In particular
$\pi^{0}\to\pi^{0}-\frac{\alpha}{2}(n_{u}-n_{d}),\qquad\eta\to\eta-\frac{\alpha}{2\sqrt{3}}(n_{u}+n_{d}-2n_{s}),\qquad\eta^{\prime}\to\eta^{\prime}-\frac{\alpha}{\sqrt{6}}(n_{u}+n_{d}+n_{s}).$ (16)
The $\cos$ terms in eq. (15) are invariant under the PQ symmetry. So the PQ shift of $\eta^{\prime}$ can be used to reabsorb the $\theta$ term, as long as $\sum_{q}n_{q}\neq 0$. A quark charged under PQ ($n_{q}\neq 0$) induces a mixing between the pseudo-scalar mesons and $a$, so that the canonical axion field is a linear combination of the pseudo-scalar mesons and $a$, weighted respectively by $f_{\pi}$ and $f_{\varphi}$.
The $4\times 4$ symmetric mass matrix $\mathcal{M}$ in the basis $\{a,\pi^{0},\eta,\eta^{\prime}\}$ is
[…]
The mixing of $a$ with the other states depends on the charges $n_{u,d,s}$ and on $f_{\pi}/f_{\varphi}$. If all quark charges are set to zero $a$ is massless. For large $f_{\varphi}\gg{\rm GeV}$, $a$ becomes weakly coupled to QCD as in the usual axion theories. In general, $f_{\varphi}$ is given by eq. (11). For charges $n_{q}\leq 1$, the (1,1) entry of the matrix can be rewritten as
$\mathcal{M}^{2}_{11}=m_{\varphi}^{2}+\lambda f_{\varphi}^{2}=m_{\sigma}^{2}-2\lambda f_{\varphi}^{2}.$ (18)
In the limit of vanishing $\lambda$, $\mathcal{M}_{11}^{2}$ is equal to $m^{2}_{\sigma}$. Depending on the size of the correction induced by mixing with the other mesons, the axion-like scalar could be slightly heavier or lighter than the radial mode. For small $f_{\varphi}$ we expect it to be slightly heavier, while the opposite occurs for large $f_{\varphi}$. As a check, the determinant of $\mathcal{M}^{2}$ is
$\det\mathcal{M}^{2}=\frac{4f_{\pi}^{2}}{3f_{\varphi}^{2}}B_{0}^{3}m_{u}m_{d}m_{s}(n_{u}+n_{d}+n_{s})^{2}m_{1}^{2}.$ (19)
It vanishes when one of the quark masses is zero, as well as if PQ is non-anomalous ($\sum_{q}n_{q}=0$) leading to a massless axion. Before analyzing the mass spectrum in various models in section 3, we consider a possibly relevant higher-order correction in the chiral Lagrangian.
2.3 Kaplan-Manohar ambiguity of the chiral Lagrangian
Going to higher orders in the chiral expansion, extra operators contribute to the chiral Lagrangian, and can be relevant for our discussion. Of particular relevance is
$\Delta\mathscr{L}=\frac{B_{0}f_{\pi}^{2}}{2}x_{\rm KM}\mathcal{O}_{\rm KM},\qquad\mathcal{O}_{\rm KM}\equiv-\frac{1}{2}\left[\operatorname{tr}(M_{q}U^{\dagger}M_{q}U^{\dagger})-\operatorname{tr}(M_{q}U^{\dagger})^{2}+\hbox{h.c.}\right].$ (20)
Such NLO term is equivalent to shifting the quark mass matrix as
$M_{q}\to M_{q}|_{\rm KM}=M_{q}+x_{\rm KM}(M_{q}^{-1})^{\dagger}\det(M_{q}^{\dagger})\det U$ (21)
in the LO operator $\operatorname{tr}(M_{q}U^{\dagger})+\hbox{h.c.}$ [10], in view of $2\det M_{q}\,\operatorname{tr}M_{q}^{-1}=(\operatorname{tr}M_{q})^{2}-\operatorname{tr}M_{q}^{2}$. The shift term has the same symmetry transformations as $M_{q}$. The presence of $\det U=e^{i\sqrt{6}\eta^{\prime}/f_{\pi}}$ guarantees that the two terms have the same transformation also under the axial U(1)A. The constant $x_{\rm KM}$ is approximately determined by lattice computations. This is known as the Kaplan-Manohar (KM) ambiguity [10], since it complicates the extraction of bare quark masses from meson data. In particular, it replaces $m_{u}\to m_{u}+x_{\rm KM}m_{d}m_{s}$. This was considered as a loophole for reviving the massless up-quark solution to the strong CP problem, before lattice calculations established that $m_{u}\neq 0$ [11]. In our case the KM ‘ambiguity’ effectively redefines the quark mass matrix into
$M_{q}|_{\rm KM}=\mathrm{diag}\left(\begin{array}[]{c}m_{u}(\sqrt{2}\varphi)^{n_{u}}+x_{\rm KM}m_{s}m_{d}(\sqrt{2}\varphi^{*})^{n_{d}+n_{s}}e^{i\sqrt{6}\eta^{\prime}}\\ m_{d}(\sqrt{2}\varphi)^{n_{d}}+x_{\rm KM}m_{u}m_{s}(\sqrt{2}\varphi^{*})^{n_{u}+n_{s}}e^{i\sqrt{6}\eta^{\prime}}\\ m_{s}(\sqrt{2}\varphi)^{n_{s}}+x_{\rm KM}m_{u}m_{d}(\sqrt{2}\varphi^{*})^{n_{u}+n_{d}}e^{i\sqrt{6}\eta^{\prime}}\end{array}\right)\,.$ (22)
The $\eta^{\prime}$ phase is necessary to reproduce the transformation under U(1)PQ and corrects the meson mass matrix introducing and extra mixing with $\eta^{\prime}$. In addition to shifting the quark masses, the KM term modifies the dependence on $\sigma$ when $n_{u}\neq n_{d}$. Its implications will be discussed in the following.
3 QCD-scale axion
The chiral Lagrangian at LO predicts that pions form a degenerate triplet (up to a tiny quark-mass-induced splitting), with the observed mass difference arising primarily from electromagnetic effects [3, 4, 5]
$\Delta_{\rm em}(m_{\pi^{0}}^{2}-m_{\pi^{\pm}}^{2})\approx-0.0013\,{\rm GeV}^{2}\sim-3\frac{\alpha_{\rm em}\Lambda^{2}}{4\pi},$ (23)
where $\Lambda\sim m_{\rho}$. Adding a QCD-scale axion generically breaks the isospin symmetry, ruining this successful prediction.
3.1 Tree-level corrections to meson masses
This class of models predicts important modifications to meson masses. To simplify formulas, we now focus on $n_{s}=0$. In the limit where the $a$ is heaviest state, i.e. $m_{\varphi}>$ GeV, we can integrate it out. This adds to the QCD squared mass matrix of neutral mesons the following additive correction
$\Delta M^{2}=B_{0}\begin{pmatrix}-\dfrac{(m_{d}n_{d}-m_{u}n_{u})^{2}}{m_{d}n_{d}^{2}+m_{u}n_{u}^{2}}&\dfrac{m_{d}^{2}n_{d}^{2}-m_{u}^{2}n_{u}^{2}}{\sqrt{3}\,(m_{d}n_{d}^{2}+m_{u}n_{u}^{2})}&\dfrac{\sqrt{\frac{2}{3}}\,(m_{d}^{2}n_{d}^{2}-m_{u}^{2}n_{u}^{2})}{m_{d}n_{d}^{2}+m_{u}n_{u}^{2}}\\[11.99998pt] \dfrac{m_{d}^{2}n_{d}^{2}-m_{u}^{2}n_{u}^{2}}{\sqrt{3}\,(m_{d}n_{d}^{2}+m_{u}n_{u}^{2})}&-\dfrac{(m_{d}n_{d}+m_{u}n_{u})^{2}}{3\,(m_{d}n_{d}^{2}+m_{u}n_{u}^{2})}&-\dfrac{\sqrt{2}\,(m_{d}n_{d}+m_{u}n_{u})^{2}}{3\,(m_{d}n_{d}^{2}+m_{u}n_{u}^{2})}\\[11.99998pt] \dfrac{\sqrt{\frac{2}{3}}\,(m_{d}^{2}n_{d}^{2}-m_{u}^{2}n_{u}^{2})}{m_{d}n_{d}^{2}+m_{u}n_{u}^{2}}&-\dfrac{\sqrt{2}\,(m_{d}n_{d}+m_{u}n_{u})^{2}}{3\,(m_{d}n_{d}^{2}+m_{u}n_{u}^{2})}&-\dfrac{2(m_{d}n_{d}+m_{u}n_{u})^{2}}{3\,(m_{d}n_{d}^{2}+m_{u}n_{u}^{2})}\end{pmatrix}$. (24)
The spectrum of charged mesons is not modified at tree level, since the new state is neutral and there is no mixing with new states. Using $m_{s}\gg m_{u,d}$, the 11 entry of the mass matrix leads to the pion mass splitting
$\frac{\Delta(m_{\pi^{0}}-m_{\pi^{\pm}})}{m_{\pi}}\approx-\frac{1}{2}\dfrac{(m_{u}n_{u}-m_{d}n_{d})^{2}}{(m_{d}n_{d}^{2}+m_{u}n_{u}^{2})(m_{u}+m_{d})}\,$, (25)
Other mesons are affected too. In particular, the $\eta$ is lighter than in the SM.
3.2 Model with axion coupled to the up quark
Murayama [1] considered the case where the up quark only is PQ-charged, $n_{u}=1$ and $n_{d}=0$. In this case one finds (neglecting electro-magnetic corrections)
$m_{\varphi}\approx 2.3\kappa_{u}{\rm GeV},\qquad m_{\pi^{0}}^{2}\approx m_{d}B_{0},\qquad m_{\pi^{\pm}}^{2}\approx B_{0}(m_{u}+m_{d}),$ (26)
that is equivalent to eq. (1). A QCD-scale $\varphi$ needs a relatively large Yukawa coupling, $\kappa_{u}\gtrsim 0.5$, not much below the perturbativity limit. The Kaplan-Manohar ambiguity helps to alleviate the problem with pion masses [1], but only mildly. It shifts the quark matrix into
$M_{q}|_{\rm KM}=\mathrm{diag}\left(\begin{array}[]{c}m_{u}\sqrt{2}\varphi+x_{\rm KM}m_{s}m_{d}e^{i\sqrt{6}\eta^{\prime}}\\ m_{d}+x_{\rm KM}m_{u}m_{s}\sqrt{2}\varphi^{*}e^{i\sqrt{6}\eta^{\prime}}\\ m_{s}+x_{\rm KM}m_{u}m_{d}\sqrt{2}\varphi^{*}e^{i\sqrt{6}\eta^{\prime}}\end{array}\right)\,.$ (27)
The KM shift terms do not modify $\langle{\varphi}\rangle$. The problematic correction to the pion mass gets suppressed if the up entry of $M_{q}|_{\rm KM}$ is dominated by the KM shift, as it is $\varphi$-independent. Assuming that the up quark mass induced by $\langle\varphi\rangle$ is a factor $R$ smaller than the measured up quark mass, we obtain
$\frac{\Delta m_{\pi^{0}}}{m_{\pi^{0}}}\approx-\frac{0.15}{R},\qquad m_{\varphi}\approx 2.3\,\kappa_{u}\,\sqrt{R}\,{\rm GeV}.$ (28)
Avoiding the pion mass problem requires $R\gtrsim 10$. This corresponds to a value of $x_{\rm KM}$ much larger than the value suggested by lattice computations, that favor a mild KM shift and disfavor reaching $m_{u}=0$, corresponding to $R=\infty$ [12, 13, 2]. Thus it seems unlikely that the KM ambiguity solves the pion mass problem.
If the model were viable, the new GeV-scale scalars should be identified with some observed hadron. Table 1 lists some candidates. Decay rates of $a$ and $\sigma$ can be estimated either from decays to quarks or to pions. Both computations are approximate and lead to similar estimates:
$\Gamma(a\to 3\pi)\approx 0.1\kappa^{2}_{u}m_{a},\qquad\Gamma(\sigma\to\pi\pi)\approx\frac{\kappa^{2}_{u}}{8\pi}m_{\sigma}.$ (29)
The relatively narrow $a$ could be identified with the observed $\eta(1295)$, as its mass is $m_{a}\sim 1.1\,{\rm GeV}$ and its width is $\Gamma_{a}\sim 25\,{\rm MeV}$ for $\kappa_{u}\sim 1/2$ (we fixed factors of 2 compared to [1]). The broad $\sigma$ can be more easily hidden inside QCD hadron resonances.
State $J^{PC}$ Mass [MeV] Width [MeV]
$\eta(547)$ $0^{-+}$ $547.9\pm 0.0$ $(1.31\pm 0.05)\times 10^{-3}$
$\eta^{\prime}(958)$ $0^{-+}$ $957.8\pm 0.1$ $0.198\pm 0.009$
$\eta(1295)$ $0^{-+}$ $1294\pm 4$ $55\pm 5$
$\eta(1405)$ $0^{-+}$ $1409\pm 2$ $51\pm 7$
$\eta(1475)$ $0^{-+}$ $1476\pm 4$ $85\pm 9$
$\eta(1760)$ $0^{-+}$ $1756\pm 9$ $96\pm 70$
$\eta(2225)$ $0^{-+}$ $2221\pm 8$ $185\pm 40$
$f_{0}(500)$ $0^{++}$ 400400550550 400400700700
$f_{0}(980)$ $0^{++}$ $990\pm 20$ 4040100100
$f_{0}(1370)$ $0^{++}$ 1200120015001500 200200500500
$f_{0}(1500)$ $0^{++}$ $1505\pm 6$ $109\pm 7$
$f_{0}(1710)$ $0^{++}$ $1723\pm 6$ $135\pm 8$
$f_{0}(2020)$ $0^{++}$ $1992\pm 16$ $442\pm 60$
$f_{0}(2200)$ $0^{++}$ $2189\pm 13$ $238\pm 50$
$f_{0}(2330)$ $0^{++}$ $2325\pm 35$ $200\pm 70$
Table 1: Isoscalar pseudoscalar ($\eta$) and scalar ($f_{0}$) resonances in QCD with their quantum numbers, masses, and total decay widths. Broad states have large systematic uncertainties.
3.3 Model with axion coupled to up and down quarks
Given that the pion mass splitting originates from the breaking of the SU(2) isospin symmetry, it is tempting to assume that both up and down quarks are PQ charged. For negligibile $\lambda$ the mass of $\sigma$ is
$m_{\varphi}=1.6\,\kappa_{d}\,{\rm GeV}\sqrt{n_{u}\frac{m_{u}}{m_{d}}+n_{d}}\,.$ (30)
[…]
$M_{q}|_{\rm KM}=\mathrm{diag}\left(\begin{array}[]{c}m_{u}\sqrt{2}\varphi+x_{\rm KM}m_{s}m_{d}\sqrt{2}\varphi^{*}e^{i\sqrt{6}\eta^{\prime}}\\ m_{d}\sqrt{2}\varphi+x_{\rm KM}m_{u}m_{s}\sqrt{2}\varphi^{*}e^{i\sqrt{6}\eta^{\prime}}\\ m_{s}+x_{\rm KM}m_{u}m_{d}(\sqrt{2}\varphi^{*})^{2}e^{i\sqrt{6}\eta^{\prime}}\end{array}\right)\,.$ (31)
A cancellation of the pion mass shift can happen for $n_{u}=2$ and $n_{d}=1$ (see eq. (25)), as data allow for $m_{d}/m_{u}=2$ within uncertainties. However $n_{u}=2$ means that the up-quark mass comes from a dimension-5 operator that gets strongly coupled at energies mildly above $m_{u}$, below the QCD scale. This would require a UV completion at that scale, with extra particles that would be problematically light.
3.4 Model with axion coupled to the strange quark
We briefly comment on the possibility that only the strange quark has a PQ charge. This would have the merit of preserving the isospin $\mathrm{SU}(2)$ symmetry of the $u,d$ sector. However one predicts
$m_{\varphi}=\left(\frac{B_{0}f_{\pi}^{2}}{2m_{s}}\right)^{1/2}\kappa_{s}\approx 0.36~\text{GeV}\,\kappa_{s}.$ (32)
for the radial mode. Upon diagonalization, the lightest pseudo-scalar turns out to be even lighter (this can be understood as the mass of the strange quark does not enter the mass matrix of neutral mesons as discussed in the next section), $m_{a}\approx 0.038~\text{GeV}$ for $\kappa_{s}=1$. Such a light axion is excluded by old collider data, and by flavour constraints, most notably by $D\to\pi a$, generated at one loop via $W$ exchange and sourced by the $a\,\bar{s}\,i\gamma_{5}s$ coupling. Moreover, the pion-mass issue is effectively shifted to the $\eta$ sector: the $\eta$ mass receives a sizeable correction, $\Delta m_{\eta}/m_{\eta}\approx+15\%$ for $\kappa_{s}=1$, which is difficult to reconcile with data.
Extra tentative models with multiple PQ scalars are discussed in appendix A. One might even consider the possibility of adding a PQ-neutral scalar just to make the $\pi^{0}$ heavier. In section 4 we discuss the general reason that prevents getting a viable model.
3.5 Loop corrections to pion masses
All results receive corrections that get large if $\kappa$ is large. Here we attempt an estimate. One-loop corrections to pion masses directly affect the dominant diagonal entries of the meson mass matrix in eq. (17). So there could be a controllable regime where the one-loop corrections cancel the too large tree-level correction to $m_{\pi^{0}}-m_{\pi^{\pm}}$, arising from off-diagonal entries in the meson mass matrix. We here heuristically derive the pion mass splitting by computing the renormalization of the meson masses in the chiral Lagrangian effective theory, including $\varphi$. In addition to isospin-symmetric vertices, it contains a $\sigma\pi^{0}\eta^{(\prime)}$ vertex that only affects the $\pi^{0}$ mass. The linear coupling of $\varphi$ to QCD mesons can be written in terms of $\varphi$-dependent meson mass matrix, obtaining the 1-loop contribution as
$\Delta M^{2}_{ij}=-(\partial_{\varphi}M^{2}_{ik})\left[\int\frac{d^{4}Q}{(2\pi)^{4}}\frac{1}{Q^{2}+m_{\varphi}^{2}}\frac{1}{Q^{2}+m^{2}}\right]_{kl}(\partial_{\varphi}M^{2}_{lj}),$ (33)
where latin indices label QCD mesons. The Euclidean integral is positive. The integral is logarithmically UV divergent. More importantly, we cannot apply the chiral Lagrangian at large $Q\gtrsim$ GeV. So we insert in the above expression a cut-off at $\Lambda\sim m_{\rho}\sim{\rm GeV}$. Assuming $m_{\varphi}\gtrsim\Lambda$ and neglecting for simplicity off-diagonal terms in the mass matrix, we find
$\Delta M^{2}_{ij}\approx-\frac{(\partial_{\varphi}M^{2}_{ik})}{16\pi^{2}m_{\varphi}^{2}}\left[\Lambda^{2}-M^{2}_{kk}\ln\frac{\Lambda^{2}+M_{kk}^{2}}{M_{kk}^{2}}\right](\partial_{\varphi}M^{2}_{kj}).$ (34)
To leading order for $n_{u}=1$ we find,
$\Delta M^{2}|_{\rm neutral}\approx-\frac{\kappa^{2}B_{0}^{2}\Lambda^{2}}{(4\pi m_{\varphi})^{2}}\begin{pmatrix}2&\frac{2}{\sqrt{3}}&2\sqrt{\frac{2}{3}}\\ 2{\sqrt{3}}&\frac{2}{3}&\frac{2\sqrt{2}}{3}\\ 2\sqrt{\frac{2}{3}}&\frac{2\sqrt{2}}{3}&\frac{4}{3}\end{pmatrix},\qquad\Delta M^{2}|_{\rm charged}\approx-\frac{\kappa^{2}B_{0}^{2}\Lambda^{2}}{(4\pi m_{\varphi})^{2}}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&0\end{pmatrix}$. (35)
This splitting, when extrapolated to $m_{\varphi}>\Lambda$ agrees, including the sign, with the result of the tree level matching eq. (24) upon identifying $\Lambda=4\pi f_{\pi}$. This however should not be considered as a new contribution to the splitting but rather as an estimate of the uncertainty.
4 Heavy axion effective chiral theory
We here focus on the regime where the new PQ scalars ($a$ and $\sigma$ in the minimal model) are heavy enough that they could be hidden among QCD resonances with GeV mass. This regime needs $n_{s}=0$ and $n_{u,d}\leq 1$ in the language of the previous section. We here exploit the (mild) separation of scales
$m_{\sigma,a}\mathrel{\raisebox{-4.30554pt}{$\overset{\gg}{\sim}$}}m_{\pi,K,\eta}$ (36)
to integrate out the PQ scalars, obtaining a modified chiral Lagrangian for the light mesons. Namely we capture the effect of the heavy field by matching the full theory of eq. (3) at $m_{\varphi}$ to a new chiral Lagrangian augmented by operators — invariant under PQ — generated by the new scalar dynamics. This allows to exploit the symmetries to constrain the form of extra effects, including higher loop effects. The low energy degrees of freedom are again the $\Pi$ fields of eq. (9). The $\eta^{\prime}$ is again kept in the effective field theory (EFT) because of its special role for the PQ symmetry. The meson interactions differ from the usual chiral Lagrangian, and even include different isospin-breaking terms. Indeed, at the matching scale, formal invariants under the full symmetry now involve two spurions:
  • The usual mass matrix of quarks, with masses now present only for PQ-neutral quarks:
    $\tilde{M}_{q}=\mathrm{diag}(m_{u}(1-n_{u}),m_{d}(1-n_{d}),m_{s}).$ (37)
  • The axion couplings to quarks, proportional to the spurion
    $I_{\rm PQ}=\mathrm{diag}(\kappa_{u}n_{u},\kappa_{d}n_{d},0).$ (38)
[…]
$I_{\rm PQ}\to I_{\rm PQ}e^{-i\alpha Q_{\rm PQ}},\qquad Q_{\rm PQ}={\rm diag}(n_{u},n_{d},0)$ (39)
under the PQ symmetry that acts on quarks as $\bar{q}_{L}q_{R}\to e^{i\alpha Q_{\rm PQ}}\bar{q}_{L}q_{R}$, and on mesons as $U\to Ue^{-i\alpha Q_{\rm PQ}}$. Integrating $\varphi$ out at the level of the QCD action generates PQ-invariant $|\bar{q}_{L}q_{R}|^{2}$ operators, while $n_{q}>1$ would lead to higher quark powers. The modified chiral EFT is constrained by the invariance under the PQ symmetry, and has the form
$\mathscr{L}_{\rm eff}=\mathscr{L}_{\rm QCD}(\tilde{M}_{q})+\sum_{i}C_{i}\mathcal{O}_{i}^{\rm PQ}.$ (40)
The first term, $\mathscr{L}_{\rm QCD}(\tilde{M}_{q})$, is the chiral Lagrangian with quark masses given by $\tilde{M}_{q}$. Like in eq. (7), it includes the term ${f_{\pi}^{2}B_{0}}\operatorname{tr}[\tilde{M}_{q}U^{\dagger}+\tilde{M}_{q}^{\dagger}U]/2$ at leading $p^{2}$ order in the chiral expansion, and terms such as $\mathcal{O}_{\rm KM}$ at higher order. The second is a sum of operators with at most two derivatives generated at the scale $m_{\varphi}$.
  • One operator first appears at $p^{4}$ order in the chiral expansion
    $\mathcal{O}_{1}^{\rm PQ}=\operatorname{tr}[I_{\rm PQ}^{\dagger}U]\operatorname{tr}[I_{\rm PQ}U^{\dagger}].$ (41)
  • Two extra operators contribute at next-to-leading $p^{6}$ order,
    $\mathcal{O}_{2}^{\rm PQ}=\operatorname{tr}[I_{\rm PQ}^{\dagger}\partial_{\mu}U]\operatorname{tr}[I_{\rm PQ}\partial^{\mu}U^{\dagger}],\qquad\mathcal{O}_{3}^{\rm PQ}=\operatorname{tr}[I_{\rm PQ}^{\dagger}(\partial_{\mu}U)(\partial^{\mu}U)^{\dagger}I_{\rm PQ}].$ (42)
In this language, the strong CP problem is solved because the operators are invariant under a global PQ rephasing of $U$, that arises when $\tilde{M}_{q}$ has at least a zero eigenvalue.
The Murayama model is now encoded in the coefficients $C_{i}$. Integrating out the PQ scalar at tree level gives
$C_{1}=\frac{1}{4}\frac{f_{\pi}^{4}B_{0}^{2}}{m_{\varphi}^{2}},\qquad C_{2}=\frac{1}{4}\frac{f_{\pi}^{4}B_{0}^{2}}{m_{\varphi}^{4}},\qquad C_{3}=0.$ (43)
At loop level one expects order unity corrections. In usual chiral perturbation theory, the coefficient $C_{1}$ would be chirally suppressed by $m_{\pi}^{2}/(4\pi f_{\pi})^{2}$, giving a suppressed contribution to pion masses. However the Murayama theory predicts a larger $C_{1}$ because ${\cal O}_{1}^{\rm PQ}$ substitutes quark masses. So the theory breaks the usual chiral expansion in powers of $p^{2}$. Unlike the LO $p^{2}$ operator, $p^{4}$ operators (including ${\cal O}_{1}^{\rm PQ}$) contribute to pion masses in a way that breaks SU(2) isospin.
4.1 Pion masses
This effective theory allows us to compute the meson masses, and see why they differ from the QCD values. The derivative operators ${\cal O}_{2,3}^{\rm PQ}$ act as a wave-function renormalization of the kinetic terms of the pions, and as such they affect the mass eigenvalues (after canonical normalization including rotation and rescaling) giving sub-leading corrections in $1/m_{\varphi}^{2}$. Using $m_{s}\gg m_{u,d}$, the masses of the neutral and charged pions are
$m_{\pi^{0}}^{2}$ $\approx$ $B_{0}\tilde{m}\bigg(1+\frac{x_{\rm KM}m_{s}}{2}\bigg)+\frac{8C_{1}\kappa_{u}\kappa_{d}n_{u}n_{d}}{f_{\pi}^{2}},$ (44)
$m_{\pi_{\pm}}^{2}$ $\approx$ $B_{0}\tilde{m}\bigg(1+\frac{x_{\rm KM}m_{s}}{2}\bigg)+\frac{2C_{1}(\kappa_{u}n_{u}+\kappa_{d}n_{d})^{2}}{f_{\pi}^{2}}.$ (45)
where $\tilde{m}\equiv m_{u}(1-n_{u})+m_{d}(1-n_{d})$. We see that $C_{1}$ breaks isospin at the level of the mass spectrum. In agreement with our previous results, $m_{\pi^{\pm}}^{2}=B_{0}(m_{u}+m_{d})$ (plus KM corrections) keeps its standard QCD value for $C_{1}$ as in eq. (43). The pion mass difference is
$m_{\pi^{0}}^{2}-m_{\pi_{\pm}}^{2}=-2\frac{C_{1}}{f_{\pi}^{2}}(\kappa_{u}n_{u}-\kappa_{d}n_{d})^{2}$ (46)
in agreement with our previous results in eq. (25), where here $n_{u,d}\leq 1$. The EFT view clarifies in a more general way why it is difficult to keep the pions degenerate. Pion masses change in a qualitative way compared to QCD because they arise from a different operator, that breaks isospin invariance. The correction to pion masses is numerically comparable to the effect of setting $m_{u}=0$ in QCD.
The Kaplan-Manohar ambiguity can help only partially, as (despite some similarity) ${\cal O}_{1}^{\rm PQ}$ differs from the Kaplan-Manohar operator in eq. (20).
Isospin SU(2) can perhaps be recovered by adding a SU(2) multiplet of extra scalars at the QCD scale, but the electrically charged components would be excluded.
Loop corrections can become more significant, since $C_{1}$ violates the chiral expansion. Nevertheless, loop corrections cannot cancel the tree-level effect, as long as one operator is dominant.
One operator dominates, so its coefficient cannot be tuned to zero in an extended theory with more GeV-scale PQ scalars. One could consider an extension of the theory where extra operators are added just to tune the pion masses to their observed values. The EFT formalism shows however that this leaves other problems, in pion interactions.
4.2 Pion interactions
In QCD, the meson mass operator is invariant under SU(2) isospin. Indeed, setting to zero the heavier mesons, its explicit form is
$\operatorname{tr}\,[M_{q}U^{\dagger}+\hbox{h.c.}]=m_{s}+(m_{u}+m_{d})\cos\pi\qquad\hbox{where}\qquad\pi^{2}\equiv\pi^{a}\pi^{a}=(\pi^{0})^{2}+2\pi^{-}\pi^{+}.$ (47)
This operator predicts a relation between the quadratic and quartic terms in the potential $V(\pi)$, namely between pion masses and couplings. As a result the 4-pion scattering amplitude at tree level in QCD has the Weinberg form
$\mathscr{A}(\pi_{a}\pi_{b}\to\pi_{c}\pi_{d})=\frac{1}{f_{\pi}^{2}}\bigg[\delta_{ab}\delta_{cd}(s-m_{\pi}^{2})+\delta_{ac}\delta_{bd}(t-m_{\pi}^{2})+\delta_{ad}\delta_{bc}(u-m_{\pi}^{2})\bigg],$ (48)
which vanishes in the soft pion limits, a feature known as Adler’s zero. The $s,t,u$ terms arise from the pion kinetic term, while the constant terms arise from the pion mass term. This term changes if the higher-order chiral operator contributes to pion masses. Indeed, in the same limit, ${\cal O}_{1}^{\rm PQ}$ gives a different isosin-breaking function of pions:
${\cal O}_{1}^{\rm PQ}=(n_{u}\kappa_{u}+n_{d}\kappa_{d})^{2}\cos^{2}\pi+(n_{u}\kappa_{u}-n_{d}\kappa_{d})^{2}\frac{(\pi^{0})^{2}}{\pi^{2}}\sin^{2}\pi.$ (49)
For some parameter values it leads to an extra local minimum, giving the cosmological signatures discussed in [14]. Tuning $n_{u}\kappa_{u}-n_{d}\kappa_{d}$ to 0 allows to restore SU(2) invariance, but with a different function of $\pi$ than in eq. (47). The quartic pion interactions, in addition to the QCD ones, are
$\mathscr{L}_{4\pi}=\mathscr{L}_{4\pi}^{\rm QCD}+B_{0}f_{\pi}^{2}\pi^{2}\frac{6\pi^{-}\pi^{+}\left(m_{d}n_{d}+m_{u}n_{u}\right)^{2}-(\pi^{0})^{2}(m_{d}^{2}n_{d}^{2}+m_{u}^{2}n_{u}^{2}-14m_{d}n_{d}m_{u}n_{u})}{24\left(m_{d}n_{d}+m_{u}n_{u}\right)}.$ (50)
Such terms contribute to pion scattering amplitudes, changing the constant term in pion scattering amplitudes from the QCD value of eq. (48). The Murayama model with $n_{d}=0$ gives a vanishing correction to scatterings involving $\pi^{0}$ only, but affects processes involving charged pions. For example, the simpler amplitudes that involve one pion mass only change into
$\mathscr{A}(\pi^{0}\pi^{0}\to\pi^{0}\pi^{0})=(4-3X_{0})\frac{m_{\pi^{0}}^{2}}{f_{\pi}^{2}},\qquad X_{0}=\left\{\begin{array}[]{ll}1&\hbox{if one $n_{q}=0$}\\ 4&\hbox{if $n_{u}=n_{d}=1$}\end{array}\right$., (51)
$\mathscr{A}(\pi^{+}\pi^{-}\to\pi^{+}\pi^{-})=\frac{s+t-2X_{\pm}m_{\pi^{\pm}}^{2}}{f_{\pi}^{2}},\qquad X_{\pm}=1+3\frac{m_{u}n_{u}+m_{d}n_{d}}{m_{u}+m_{d}}.$ (52)
Here $X_{0,\pm}$ are the relevant $V^{\prime\prime\prime\prime}/V^{\prime\prime}$ terms, normalized such that the standard QCD result corresponds to $X_{0,\pm}=1$. The difference in the pion cross section is of order unity in the non-relativistic limit $s\to 4m_{\pi}^{2}$, $t\to 0$. Such cross sections are measured with few $\%$ precision to agree with the QCD values, see e.g. [15].
5 Light invisible axion
A heavy axion resembling a QCD meson would be a new solution to the strong CP problem, but we could not solve its problems. We conclude by considering the same theory in the limiting regime where the axion is much lighter than the QCD scale, so that the EFT of section 4 is no longer applicable. As $f_{\varphi}$ becomes large, the pseudo-scalar component $a$ (and its scalar companion $\sigma$) become lighter, producing an invisible axion scenario with the special property that QCD itself breaks the PQ symmetry, leading to a non-standard cosmological history. In this region of the parameter space the masses of the two scalars read (in the following $n_{q}\leq 1$)
$m_{a}^{2}=\frac{m_{u}m_{d}}{(m_{u}+m_{d})^{2}}\frac{m_{\pi}^{2}f_{\pi}^{2}}{f_{\varphi}^{2}}\left(\sum_{q}n_{q}\right)^{2},\quad\quad m_{\sigma}^{2}=\frac{m_{\pi}^{2}f_{\pi}^{2}}{f_{\varphi}^{2}}\frac{\sum_{q}m_{q}n_{q}}{m_{u}+m_{d}}+2\lambda f^{2}_{\varphi},$ (53)
where the former can be obtained from diagonalizing the pseudoscalar mass matrix in the limit $f_{\varphi}\gg f_{\pi}$. In this regime the canonically normalized axion field can be identified with $a$. One may then use the standard axion mass relation [16], expressed in terms of the axion decay constant $f_{a}$, to determine the matching condition
$f_{a}\equiv\frac{f_{\varphi}}{\sum_{q}n_{q}}\,.$ (54)
A light axion is subject to the usual astrophysical constraints $f_{\varphi}\gtrsim 10^{8}$ GeV [6]. Reproducing the desired quark masses needs small Yukawa couplings $\kappa_{q}$,
$\kappa_{u}=\sqrt{2}\frac{m_{u}}{f_{\varphi}}=3.1~10^{-11}\frac{10^{8}\,{\rm GeV}}{f_{\varphi}},$ (55)
in the model where $m_{u}$ only arises from PQ breaking.
The assumption that PQ breaking is triggered by the QCD condensate has novel cosmological consequences, but also a novel problem. Getting $f_{\varphi}\gg f_{\pi}$ from the small QCD-induced PQ breaking needs a small scalar quartic. So both scalars $a$ and $\sigma$ become extremely light, in a comparable way. This can be seen applying the inequality of eq. (12) to the mass relations of eq. (53), obtaining
$m^{2}_{\sigma}=\frac{m_{u}+m_{d}}{m_{d}}m^{2}_{a}+2\lambda f^{2}_{\varphi}\lesssim\frac{m_{u}+m_{d}}{m_{d}}m^{2}_{a}+10^{-4}\,\text{GeV}^{2}\left(\frac{\text{GeV}}{f_{\varphi}}\right)^{2},$ (56)
where we again focused on the model where only $m_{u}$ arises from PQ breaking. This eventually implies the upper bound $m_{\sigma}\lesssim 2m_{a}$.
This is problematic due to stringent fifth-force constraints from the exchange of the $\sigma$ scalar. The effective scalar-nucleon coupling $g_{\sigma N}$, defined by the interaction term $\mathscr{L}\supset-g_{\sigma N}\,\sigma_{\rm can}\,\bar{N}N$, is given in the isospin limit by [17]
$g_{\sigma N}\approx\sum_{q}\frac{\sigma_{q}}{m_{q}}\,y_{q},$ (57)
where $y_{q}$ is the scalar-quark coupling defined through $\mathscr{L}\supset-y_{q}\,\sigma_{\rm can}\,\bar{q}q$, and the nucleon sigma terms are approximately $\sigma_{u}\approx 16~\text{MeV}$, $\sigma_{d}\approx 22~\text{MeV}$, and $\sigma_{s}\approx 50~\text{MeV}$. Using eq. (3), $y_{q}=n_{q}\,m_{q}/f_{\varphi}$, we obtain
$g_{\sigma N}\approx\sum_{q}n_{q}\,\frac{\sigma_{q}}{f_{\varphi}}\approx 1.6\times 10^{-12}\left(\frac{10^{10}\,\text{GeV}}{f_{\varphi}}\right),$ (58)
where the last benchmark value corresponds again to the $\{n_{u},n_{d},n_{s}\}=\{1,0,0\}$ model.
The scalar coupling to nucleons induces a Yukawa-like non-relativistic potential (not spin-suppressed). For sufficiently light scalars (in particular in the sub-eV range), this new force can compete with gravity and is therefore tightly constrained by tests of the inverse-square law and searches for equivalence-principle violation (see e.g. [18]). As a representative benchmark, current limits translate into $g_{\sigma N}\lesssim 10^{-13}$ for $m_{\sigma}\approx 1~\text{eV}$ (near the edge of astrophysical constraints), tightening to $g_{\sigma N}\lesssim 10^{-24}$ for $m_{\sigma}\approx 10^{-16}~\text{eV}$, basically ruling out the whole parameter space in the light invisible axion limit. Similar results are obtained in models with $n_{q}=2$, in which the PQ symmetry breaking is triggered by the QCD quark condensate by turning the sign of the $|\varphi|^{2}$ mass term. Also in this case the radial mode cannot be decoupled from the axion and fifth-force constraints exclude this scenario.
Therefore the only viable option appears to add a larger source of spontaneous PQ breaking (such as a larger $m_{\varphi}^{2}<0$), recovering a usual DFSZ-like scenario.
6 Conclusions
We critically examined the possibility proposed in [1] that the axion might be hidden within hadronic resonances. The novel idea is to introduce a complex scalar field charged under a PQ symmetry, whose breaking is triggered by the QCD chiral condensate. In this way, the up-quark mass term is dynamically generated after confinement, reviving the logic of the $m_{u}=0$ solution to the strong CP problem. If realized in nature, this mechanism would imply that the new scalar degrees of freedom are already observed among known hadrons, with the $\eta(1295)$ being a plausible candidate. However the model predicts an order unity splitting among charged and neutral pions [1]. We confirm this problem, fixing order unity factors. Avoiding this problem relying on the Kaplan-Manohar ambiguity to set $m_{u}\ll m_{d}$ is almost as excluded as using it to more simply set $m_{u}=0$.
Motivated by the hope of avoiding this phenomenological exclusion, we considered uncertainties in pion masses and explored several extensions of the minimal setup. Nevertheless, despite some partial success, we find that this entire class of models is strongly constrained, and in practice excluded, by low-energy mesonic observables. The underlying difficulty can be formulated as a general obstruction. If the new scalar degree(s) of freedom are heavy enough to be hidden among hadronic resonances, they can be integrated out, leading to a modified chiral effective theory for the mesons. The presence of the PQ symmetry introduces a new spurion that explicitly breaks isospin. As a consequence, one novel dominant operator appears in the chiral Lagrangian ($\mathcal{O}_{1}^{\rm PQ}$ in eq. (41)), with a coefficient that must make its effects comparable to those of usual chiral operator arising from quark masses. This alternative PQ-symmetric operator is problematic. At quadratic order in pions, it breaks the mass degeneracy among them. At quartic order this operator induces order-one deviations from the well-tested low-energy pion-pion scattering amplitudes.
While one might attempt to cure the pion mass problem by introducing additional operators and tuning parameters, such modifications leave corrections to meson interactions. The core issue is therefore structural: the proposed mechanism requires a non-standard chiral effective action, with new spurions that break isospin.
Finally, we have considered the opposite limit in which the PQ breaking from QCD leads to a large axion decay constant $f_{a}\gg{\rm GeV}$ approaching the usual regime of a light invisible axion. To achieve this, the PQ scalar needs to have a small quartic. As a result its radial mode is almost as light as the axion. The resulting light scalar mediates a Yukawa-type force between nucleons, which is subject to stringent fifth-force constraints. These bounds exclude large portions of the parameter space relevant for solving the strong CP problem.
Acknowledgements
LDL is supported by the European Union – Next Generation EU and by the Italian Ministry of University and Research (MUR) via the PRIN 2022 project n. 2022K4B58X – AxionOrigins. The work of MR and AT is supported by the Italian Ministry of University and Research (MUR) via the PRIN 2022 project n. 20228WHTYC (CUP:I53C24002320006). AVT is funded by the European Union, NextGenerationEU, National Recovery and Resilience Plan (mission 4, component 2) under the project MODIPAC: Modular Invariance in Particle Physics and Cosmology (CUP C93C24004940006). AS gratefully acknowledges the absence of financial support.
Appendix A Renormalizable models with more PQ scalars
As an attempt to tune the pion masses, one can consider models with more than one GeV-scale PQ scalars and thereby more free parameters. We here consider two scalars, $\varphi_{u}$ and $\varphi_{d}$, that give mass to the quark $u$ and $d$ respectively. The Lagrangian for canonically normalized fields reads
$\mathscr{L}\supset|\partial\varphi_{u}|^{2}+|\partial\varphi_{d}|^{2}-m_{\varphi_{u}}^{2}|\varphi_{u}|^{2}-m_{\varphi_{d}}^{2}|\varphi_{d}|^{2}+(\kappa_{u}\varphi_{u}\overline{u}_{L}u_{R}+\kappa_{d}\varphi_{d}\overline{d}_{L}d_{R}+\text{h.c.})-V_{\rm PQ}(\varphi_{u},\varphi_{d})\,.$ (59)
The potential $V_{\rm PQ}(\varphi_{u},\varphi_{d})$ preserves one combination of the two U(1)’s acting on the two scalars, identified as the PQ symmetry. The potential contains only one term of the type $(\varphi_{u}^{n_{d}}\varphi_{d}^{n_{u}}+\text{h.c.})$ or $(\varphi_{u}^{n_{d}}{\varphi_{d}^{*}}^{n_{u}}+\text{h.c.})$. These terms effectively realize the following charge assignments $(n_{u},n_{d})=\{(1,1),(2,\pm 1),(\pm 1,2),(3,\pm 1),(\pm 1,3)\}$, if only one of such terms is present at a time.
Writing the vacuum expectation values as $\langle\varphi_{u,d}\rangle\equiv f_{u,d}/\sqrt{2}$, the spontaneous breaking of the two U(1)’s together with their diagonal breaking via $V_{\rm PQ}$ allows us to identify the scale $f_{\varphi}$, appearing in eq. (3), as
$\varphi_{u,d}=\frac{f_{u,d}+\sigma_{u,d}}{\sqrt{2}}e^{ia_{u,d}/f_{u,d}}\,,\quad\quad f_{\varphi}^{2}\equiv f_{u}^{2}n_{d}^{2}+f_{d}^{2}n_{u}^{2}.$ (60)
The $V_{\rm PQ}$ term allows us to identify the new pseudoscalar heavy state $A$, corresponding to the linear combination $n_{d}a_{u}/f_{u}\pm n_{d}a_{d}/f_{d}$, where the $\pm$ refers to terms without and with the complex conjugation. The axion $a$ is the orthogonal combination. Diagonalization brings the scalars to the form
$\varphi_{u}=\frac{f_{u}+\sigma_{u}}{\sqrt{2}}\exp\bigg(in_{u}\frac{a}{f_{\varphi}}+i\frac{f_{d}}{f_{\varphi}}n_{d}\frac{A}{f_{\varphi}}\bigg)\,,\qquad\varphi_{d}=\frac{f_{d}+\sigma_{d}}{\sqrt{2}}\exp\bigg(\pm(in_{d}\frac{a}{f_{\varphi}}-i\frac{f_{u}n_{u}}{f_{\varphi}}\frac{A}{f_{\varphi}})\bigg)\,$, (61)
where now both $a$ and $A$ have canonically normalized kinetic terms and $\pm$ corresponds to the term with and without the complex conjugation. This parametrization correctly identifies the charges of the SM quarks and the action of PQ on the Goldstone boson $a$. Indeed, in eq. (59) the axion $a$ now only appears in the Yukawa terms, while $A$ also appears in $V_{\rm PQ}$. The mass of $A$ is controlled by the size of $V_{\rm PQ}$. We focus on PQ charges $(n_{u},n_{d})=(1,1)$, such that $V_{\rm PQ}$ is simply a mass term:
$V_{\rm PQ}(\varphi_{u},\varphi_{d})\equiv\mu_{ud}^{2}\,\varphi_{u}\varphi_{d}^{*}+\text{h.c.}.$ (62)
This interaction breaks the rotation of up and down quark to the diagonal combination where the quarks have the same charge. In this case, we find
$\langle\varphi_{u}\rangle\equiv\frac{f_{u}}{\sqrt{2}}=\frac{B_{0}f_{\pi}^{2}}{2}\,\frac{\kappa_{u}m_{\varphi_{d}}^{2}-\kappa_{d}\mu_{ud}^{2}}{m_{\varphi_{u}}^{2}m_{\varphi_{d}}^{2}-\mu_{ud}^{4}}\qquad\text{and}\qquad\langle\varphi_{d}\rangle\equiv\frac{f_{d}}{\sqrt{2}}=\frac{B_{0}f_{\pi}^{2}}{2}\,\frac{\kappa_{d}m_{\varphi_{u}}^{2}-\kappa_{u}\mu_{ud}^{2}}{m_{\varphi_{u}}^{2}m_{\varphi_{d}}^{2}-\mu_{ud}^{4}}\,.$ (63)
Consequently,
$m_{\varphi_{u}}^{2}$ $=\frac{B_{0}f_{\pi}^{2}\kappa_{u}}{\sqrt{2}f_{u}}-\frac{f_{d}}{f_{u}}\mu_{ud}^{2}=\frac{B_{0}f_{\pi}^{2}}{2}\,\frac{\kappa_{u}^{2}}{m_{u}}-\frac{\kappa_{u}}{\kappa_{d}}\,\frac{m_{d}}{m_{u}}\,\mu_{ud}^{2}\,$, (64)
$m_{\varphi_{d}}^{2}$ $=\frac{B_{0}f_{\pi}^{2}\kappa_{d}}{\sqrt{2}f_{d}}-\frac{f_{u}}{f_{d}}\mu_{ud}^{2}=\frac{B_{0}f_{\pi}^{2}}{2}\,\frac{\kappa_{d}^{2}}{m_{d}}-\frac{\kappa_{d}}{\kappa_{u}}\,\frac{m_{u}}{m_{d}}\,\mu_{ud}^{2}\,$, (65)
where in the last equalities we have taken into account that
$m_{u}=\frac{\kappa_{u}f_{u}}{\sqrt{2}}\qquad\text{and}\qquad m_{d}=\frac{\kappa_{d}f_{d}}{\sqrt{2}}\,.$ (66)
Expanding the scalars in components as
$\varphi_{u}=\frac{f_{u}+\sigma_{u}}{\sqrt{2}}e^{i\frac{a_{u}}{f_{u}}}\qquad\text{and}\qquad\varphi_{d}=\frac{f_{d}+\sigma_{d}}{\sqrt{2}}e^{i\frac{a_{d}}{f_{d}}}\,$, (67)
we see that the $\mu$-term generates a contribution to the potential of $a_{u,d}$ of the form
$\mu_{ud}^{2}f_{u}f_{d}\cos\left(\frac{a_{u}}{f_{u}}-\frac{a_{d}}{f_{d}}\right)\,.$ (68)
Adding it to the chiral Lagrangian leads to the following squared mass matrix of pseudo-scalars:
$\mathcal{M}^{2}=\begin{pmatrix}m_{\varphi_{u}}^{2}&\mu_{ud}^{2}&-\frac{B_{0}f_{\pi}\kappa_{u}}{\sqrt{2}}&-\frac{B_{0}f_{\pi}\kappa_{u}}{\sqrt{6}}&-\frac{B_{0}f_{\pi}\kappa_{u}}{\sqrt{3}}\\[5.69046pt] -&m_{\varphi_{d}}^{2}&\frac{B_{0}f_{\pi}\kappa_{d}}{\sqrt{2}}&-\frac{B_{0}f_{\pi}\kappa_{d}}{\sqrt{6}}&-\frac{B_{0}f_{\pi}\kappa_{d}}{\sqrt{3}}\\[5.69046pt] -&-&B_{0}(m_{u}+m_{d})&\frac{B_{0}}{\sqrt{3}}(m_{u}-m_{d})&\sqrt{\frac{2}{3}}B_{0}(m_{u}-m_{d})\\[5.69046pt] -&-&-&\frac{B_{0}}{3}(m_{u}+m_{d}+4m_{s})&\frac{\sqrt{2}}{3}B_{0}(m_{u}+m_{d}-2m_{s})\\[5.69046pt] -&-&-&-&\frac{2}{3}B_{0}(m_{u}+m_{d}+m_{s})+m_{1}^{2}\end{pmatrix}$. (69)
Its determinant reads
$\det\mathcal{M}^{2}=-\frac{8}{3}B_{0}^{3}f_{\pi}^{2}m_{1}^{2}m_{s}\mu_{ud}^{2}\kappa_{u}\kappa_{d}\,.$ (70)
If $\kappa_{u,d}>0$, one needs $\mu_{ud}^{2}<0$ to have all pseudo-scalar masses squared positive.
We performed a scan of the parameter space of this model, but we do not find values that reproduce the observed spectrum of $\pi^{0}$, $\eta$, $\eta^{\prime}$ while all extra scalars have heavy enough GeV masses. Consistently with the effective theory argument of section 4, fixing the pion mass problem needs extra lighter scalars, that cannot be hidden among QCD resonances. A similar issue is found studying more general models.